A share of BAC common stock has just paid a dividend of $1.00. The market return is 12% and the beta is 1.5. The three month T-bill rate is 4%. The expected long-run growth rate for this stock is 8 percent.

(Keep your answer to only two decimals)
a. What is the required return for the stock ? (hint: CAPM) (Example of the answer format: 55.55% )
b.  What is the stock price?  (Example of the answer format: $55.55))
Answer #1

a.required return= risk-free rate +Beta*(market rate- risk-free rate )

=4+1.5*(12-4)

=16%

b.Current price=D1/(Required return-Growth rate)

=(1*1.08)/(0.16-0.08)

=(1.08/0.08)

=$13.50